A ball is thrown upward from the ground with an initial speed of 18.2 m/s; at the same instant a ball is dropped from rest from a building 12 m high. After how long will the balls be at the same height?
1
Expert's answer
2011-08-31T06:47:16-0400
Let's write the equations of motion for both objects:
1. H1= H10 + v10t - gt2/2, H10=0, v10 = 18.2 m/s, g = 9.8 m/s2
2. H2 = H20 + v20t - gt2/2, H20=12, v20 = 0 m/s, g = 9.8 m/s2


Thus
H1 = H2
v10t - gt2/2 = H20 - gt2/2

v10t = H20

t = H20/v10 = 12/18.2 = 0.66 s